Can someone please help me? Thanks!

Problem 7

a. Calculate the number of C atoms in 0.751 mole of C.

b. Calculate the number of SO2 molecules in 4.36 moles of SO2.

c. Calculate the moles of Fe in 5.66x1022 atoms of Fe.

d. Calculate the moles of C2H6O in 2.07x1024 molecules of C2H6O.
Answer #1

a)
use:
number of atoms = number of moles * Avogadro’s number
= 0.751 * 6.022*10^23 atoms
= 4.523*10^23 atoms
Answer: 4.52*10^23 atoms

b)
use:
number of molecules = number of mol * Avogadro’s number
number of molecules = 4.36 * 6.022*10^23 molecules
number of molecules = 2.626*10^24 molecules
Answer: 2.63*10^24 molecules

c)
use:
number of mol = number of atoms / Avogadro’s number
number of mol = 5.66*10^22 / 6.022*10^23 mol
number of mol = 9.399*10^-2 mol
Answer: 9.40*10^-2 mol

d)
use:
number of mol = number of molecules / Avogadro’s number
number of mol = 2.07*10^24 / 6.022*10^23 mol
number of mol = 3.437 mol
Answer: 3.44 mol